Express the limit as a definite integral. n lim Σ 1P10k1 TCK AXk, where P is a partition of [6, 12] 6 OA. 7x6 dx 12 n B. 7xFind the derivative. to y = = S cos Vt dt 0 O A. cos (x3) O B. sin (x3) OC. 6x5 OD. cos (x3) - 1 cos (x3)Solve the initial value problem. dy = x(2+x2)), y(0) = 0 dx 1 O A. y= 3 (2+x2) OB. (2+x2)*-2 OC. y= 1 (2+x2) OD. y=1(2+x2)4-4
Answer #1

flag du 1 =1 We know that A= lim E f (xi) AXi = 1P10 given that lim ARK 110110 k=1 y= Sasse de find the derivative dy d . dnElekt 6) x = O AI QUAD CAMERA OO REDMI NOTE 8 hp up. hp 3 x/2+x²) X both side integration Say= 2(x + 2) du s let x + 9 = M z